# A New Application of the Gibbons-Werner Method: Bound Orbits of Massive Particles in Stationary Spacetimes

### v1

@@ Abstract
I Introduction
Einstein’s general relativity is often summarized by the statement: ”Spacetime tells matter how to move; matter tells spacetime how to curve”. A natural implication of the first half of this statement is that the motion of particles—whether massless or massive—encodes information about the background spacetimes. In relativistic astrophysics, extracting observable quantities from particle trajectories in curved spacetimes is therefore a topic of central importance.
@@ I Introduction
This paper is organized as follows. Sec. II provides a brief review of the generalized GW method for SAS spacetimes. In Sec. III, we extend the GW method to bound massive particles in SAS spacetimes and discuss the corresponding observational implications. Sec. IV demonstrates the computational technique with the Kerr spacetime as an example. Finally, Sec. V presents the conclusion. Throughout this paper, the spacetime signature (,+,+,+-,+,+,+) and the geometric units $G=c=1$ are adopted.
II Generalized GW method for SAS spacetimes
II.1 JMRF metric
@@ II.2 Generalized GW method
III Extending GW method to bound massive particles in SAS spacetimes
@@
III.1 Constraints on JMRF metric from bound orbits
@@
III.2 Deflection angle of bound orbits
@@
For the segment $\overset{\curvearrowright}{AM}$, the auxiliary circular arc $C$ intersects with the outgoing radial curves $\overrightarrow{LA}$ and $\overrightarrow{LM}$ at $A_{c}$ and $M_{c}$, respectively. Then we obtain a quadrilateral region $D_{1}=^{M_{c}}_{M}\square^{A_{c}}_{A}$. Similar to the scenario in Sec. II.2, applying the GBT to $D_{1}$ brings about
@@
$\iint_{D_{2}}K\mathrm{d}S+\int_{\overset{\curvearrowright}{M_{c}Q_{1}}}\kappa\mathrm{d}l+\int_{\overset{\curvearrowright}{Q_{1}M}}\kappa\mathrm{d}l+\int_{\overrightarrow{MM_{c}}}\kappa\mathrm{d}l$
$+\eta_{M}+\eta_{M_{c}}+\eta_{Q_{1}}=2\pi\xi(D_{2}),$ (21)
$\iint_{D_{3}}K\mathrm{d}S+\int_{\overset{\curvearrowright}{Q_{1}\gamma Q_{2}}}\kappa\mathrm{d}l+\int_{\overset{\curvearrowright}{Q_{2}CQ_{1}}}\kappa\mathrm{d}l$
$+\eta_{Q_{1}}+\eta_{Q_{2}}=2\pi\chi(D_{3}),$ (22)
$\iint_{D_{4}}K\mathrm{d}S+\int_{\overset{\curvearrowright}{Q_{2}N_{c}}}\kappa\mathrm{d}l+\int_{\overrightarrow{N_{c}N}}\kappa\mathrm{d}l+\int_{\overset{\curvearrowright}{NQ_{2}}}\kappa\mathrm{d}l$
$+\eta_{Q_{2}}+\eta_{N_{c}}+\eta_{N}=2\pi\chi(D_{4}).$ (23)
@@ III.3 Correspondence to observations
IV Calculation in Kerr spacetime
@@ IV.1 Equation of motion
$\mathcal{B}=$ $2\left(3+e^{2}\right)M^{2}-2a^{2}-2Mp-\frac{2a^{2}\left(3+e^{2}\right)M}{p}$
$-\frac{4a^{2}\left(e^{2}-1\right)M}{p},$
@@
The ”++” of the operator $\pm$” in Eq. (44) corresponds to the retrograde orbit, and ”-” the prograde orbit Bini et al. (2016). For the prograde orbit, we set $\mathcal{L}>0$ and $a>0$, thus
@@ IV.3 Deflection angle
displays $\delta_{BA}$, $\phi_{BA}$, and $\Psi_{BA}$ as functions of $\chi_{B}$ (the decomposition $\delta_{BA}=\Psi_{BA}+\phi_{BA}$ is discussed in Sec. II.2). Without loss of generality, we set $\xi_{A}=0$, namely, point $A$ is assumed to lie at a pericenter. The behavior of $\delta_{BA}$ is dominated by $\phi_{BA}$, while $\Psi_{BA}$ contributes a periodic modulation. $\phi_{BA}$ increases monotonically as point $B$ moves along the orbit, it slightly exceeds $2\pi$ when $B$ reaches the adjacent pericenter ($\chi_{B}=2\pi$), corresponding to the pericenter advance angle. $\Psi_{BA}$ varies periodically with a period of $2\pi$ and becomes zero when point $B$ reaches a pericenter or an apocenter, since the angle between the velocity and the outgoing radial direction is $\pi/2$ at both locations. Moreover, $\Psi_{BA}$ is more sensitive near the apocenter than near the pericenter. Fig. 6

### v2

# Applying the Gibbons-Werner method to bound orbits of massive particles in stationary spacetimes

@@ Abstract
journal: Eur. Phys. J. C
1 Introduction
Einstein’s general relativity is often summarized by the statement: "Spacetime tells matter how to move; matter tells spacetime how to curve". A natural implication of the first half of this statement is that the motion of particles—whether massless or massive—encodes information about the background spacetimes. In relativistic astrophysics, extracting observable quantities from particle trajectories in curved spacetimes is therefore a topic of central importance.
@@ 1 Introduction
This paper is organized as follows. Sec. 2 provides a brief review of the generalized GW method for SAS spacetimes. In Sec. 3, we extend the GW method to bound massive particles in SAS spacetimes and discuss the corresponding observational implications. Sec. 4 demonstrates the computational technique with the Kerr spacetime as an example. Finally, Sec. 5 presents the conclusion. Throughout this paper, the spacetime signature (,+,+,+-,+,+,+) and the geometric units $G=c=1$ are adopted.
2 Generalized GW method for SAS spacetimes
2.1 JMRF metric
@@ 2.2 Generalized GW method
3 Extending GW method to bound massive particles in SAS spacetimes
@@
3.1 Constraints on JMRF metric from bound orbits
@@
3.2 Deflection angle of bound orbits
@@
For the segment $\overset{\curvearrowright}{AM}$, the auxiliary circular arc $C$ intersects with the outgoing radial curves $\overrightarrow{LA}$ and $\overrightarrow{LM}$ at $A_{c}$ and $M_{c}$, respectively. Then we obtain a quadrilateral region $D_{1}=^{M_{c}}_{M}\square^{A_{c}}_{A}$. Similar to the scenario in Sec. 2.2, applying the GBT to $D_{1}$ brings about
@@
$\iint_{D_{2}}K\mathrm{d}S+\int_{\overset{\curvearrowright}{M_{c}Q_{1}}}\kappa\mathrm{d}l+\int_{\overset{\curvearrowright}{Q_{1}M}}\kappa\mathrm{d}l+\int_{\overrightarrow{MM_{c}}}\kappa\mathrm{d}l$ (21)
$+\eta_{M}+\eta_{M_{c}}+\eta_{Q_{1}}=2\pi\xi(D_{2}),$
$\iint_{D_{3}}K\mathrm{d}S+\int_{\overset{\curvearrowright}{Q_{1}\gamma Q_{2}}}\kappa\mathrm{d}l+\int_{\overset{\curvearrowright}{Q_{2}CQ_{1}}}\kappa\mathrm{d}l$ (22)
$+\eta_{Q_{1}}+\eta_{Q_{2}}=2\pi\chi(D_{3}),$
$\iint_{D_{4}}K\mathrm{d}S+\int_{\overset{\curvearrowright}{Q_{2}N_{c}}}\kappa\mathrm{d}l+\int_{\overrightarrow{N_{c}N}}\kappa\mathrm{d}l+\int_{\overset{\curvearrowright}{NQ_{2}}}\kappa\mathrm{d}l$ (23)
$+\eta_{Q_{2}}+\eta_{N_{c}}+\eta_{N}=2\pi\chi(D_{4}).$
@@ 3.3 Correspondence to observations
4 Calculation in Kerr spacetime
@@ 4.1 Equation of motion
$\mathcal{B}=$ $2\left(3+e^{2}\right)M^{2}-2a^{2}-2Mp$
$-\frac{4a^{2}\left(e^{2}-1\right)M}{p}-\frac{2a^{2}\left(3+e^{2}\right)M}{p},$
@@
The "++" of the operator $\pm"$ in Eq. (44) corresponds to the retrograde orbit, and "-" the prograde orbit Bini et al. (2016). For the prograde orbit, we set $\mathcal{L}>0$ and $a>0$, thus
@@ 4.3 Deflection angle
displays $\delta_{BA}$, $\phi_{BA}$, and $\Psi_{BA}$ as functions of $\chi_{B}$ (the decomposition $\delta_{BA}=\Psi_{BA}+\phi_{BA}$ is discussed in Sec. 2.2). Without loss of generality, we set $\xi_{A}=0$, namely, point $A$ is assumed to lie at a pericenter. The behavior of $\delta_{BA}$ is dominated by $\phi_{BA}$, while $\Psi_{BA}$ contributes a periodic modulation. $\phi_{BA}$ increases monotonically as point $B$ moves along the orbit, it slightly exceeds $2\pi$ when $B$ reaches the adjacent pericenter ($\chi_{B}=2\pi$), corresponding to the pericenter advance angle. $\Psi_{BA}$ varies periodically with a period of $2\pi$ and becomes zero when point $B$ reaches a pericenter or an apocenter, since the angle between the velocity and the outgoing radial direction is $\pi/2$ at both locations. Moreover, $\Psi_{BA}$ is more sensitive near the apocenter than near the pericenter. Fig. 6
@@ 5 Conclusion
Acknowledgments
This work was supported in part by Natural Science Foundation of Hunan Province for Youths Grant No. 2024JJ6210 and 2024JJ6211, the National Natural Science Foundation of China Grant No. 12405054, 12375045, 12205093, and 12405053, and the Science Research Fund of Hunan Provincial Education Department No. 24B0476, 21A0297 and 24C0229.